Use the equation 3x + 5y = 0
1. To Rewrite the equation above in Slope-Intercept format, (solve for y).
2. Then find the slop and y intercept..
3x + 5y = 0
5y = -3x
y= -3/5 * x and b=0
How would I then :
3. Find the Slope and y intercept.

Thank you, I greatly appreciate the help.

Answer by stanbon(75887)

Use the equation 3x + 5y = 0
1. To Rewrite the equation above in Slope-Intercept format, (solve for y).
2. Then find the slop and y intercept..
3x + 5y = 0
5y = -3x
y= -3/5 * x and b=0
How would I then :
3. Find the Slope and y intercept.
The y-intercept is f(0) = (-3/5)0 = 0
y-intercept:: 0 or (0,0)
--------------------------------
The Slope is the change in "y" when x changes by "1"::
Remember: when x = 0, y = 0
Find f(1) = (-3/5)*1 = -3/5
The change is "-3/5"
So the slope is -3/5
-----------------------
In general, if y = mx + b, m is the slope and b is the y-intercept
------------------
Cheers,
Stan H.
-----------